OpenStudy (anonymous):

Solve the system. y = x + 2 and y = -x + 0 A. (1, -1) B. (-1, 1) C. (1, 3) D. (-1, 3) E. no solution

OpenStudy (anonymous):

B

OpenStudy (anonymous):

y=x+2.............1 y=-x+0............2 sub 2 in 1 x+2=-x+0 2x=-2 x=-1 sub this value of x in any equaiton y=-1+2 = 1 (-1,1)